### v4

# Group-Adaptive Threshold Optimization for Robust AI-Generated Text Detection

Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, one universal threshold could fail to account for distributional variations by subgroups. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text, and more positive classifications of neurotic writing styles among long texts. These discrepancies can lead to misclassifications that disproportionately affect certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization for probabilistic AI-text detectors. We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy. In experiments with nine AI text classifiers on three datasets, FairOPT decreases overall balanced error rate (BER) discrepancy by 12% while minimally sacrificing accuracy by 0.003%. Our framework paves the way for more robust classification in AI-generated content detection via post-processing. We release our data, code, and project information at URL.
1 Introduction
The copious use of large language models (LLMs) facilitated widespread dissemination of AI-generated text [1, 2] and has raised concerns around the dissemination of misinformation [3], the degradation of publication standards [4], cybersecurity threats [5], and breaches of academic integrity [6]. In response to the anticipated risks, AI detection techniques have been proposed [1, 7]. Typically, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities) of the given text against AI-generated text, and use a universal threshold (e.g., $\theta>0.5$[8] for a binary classification.
Figure 1: We find that using one universal probability threshold (often $\theta>0.5$) to classify AI and human-generated text can fail to account for subgroup-specific distributional variations, such as text length or writing style, leading to higher errors for certain demographics. We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which reduces discrepancy in AI text detection as shown in Figure 7.
Figure 2: Applying adaptive thresholds to different probability distributions by subgroup in the test dataset, as seen through cumulative density functions (CDFs). We partition texts by three length categories and five personality traits as noted in Table 1. Observe that a static classification threshold ($0.5$, in black) and a single optimized threshold using AUROC-based method (in gray) do not account for subgroup-specific distributional variations. We used fifteen thresholds that have been generated using FairOPT to address the varying probabilistic distribution from AI text classifiers. The training is based on the RoBERTa-large detector. Details about the thresholds are in Appendix E.
However, a single universal threshold for AI content classification neglects the heterogeneity of probability distributions by textual characteristics, such as length and writing style. To put in another way, one fixed threshold does not consider the variation of the probability from various stylistic features. For example, when using a fixed threshold, a writing style characterized by long neurotic writing could be more likely to be identified as AI-generated than one characterized by agreeableness, resulting in disproportionate misclassifications, affecting certain groups. Analogous disparities have been documented in other classification tasks, regarding race [9, 10] and gender [11, 12, 10].
To overcome these limitations, we introduce FairOPT, a group-adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1). Our method balances performance (e.g., accuracy (ACC), F1, precision) and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds. Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces BER disparity across subgroups while incurring only a marginal tradeoff in overall performance.
To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving a 12% reduction in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
2 Related Work
AI text detectors:
AI text detectors have been developed to tackle issues related to misinformation, integrity, and security. RADAR utilizes adversarial learning techniques [13], GLTR employs statistical metrics, including entropy, probability distribution over the model’s vocabulary for each next token (i.e. the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [14], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [15], and RoBERTa-based detectors fine-tune pretrained language models [16]. We employed probabilistic classifiers for an AI-likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
Static thresholding for classification:
The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [8]. This default rule is commonly used in many fields like photogrammetry [17], ecology [18, 19], and machine learning [20]. Nevertheless, as highlighted by Freeman and Moisen [8], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
Universal thresholding with optimization:
Threshold optimization techniques have been formulated to learn an optimal threshold across the dataset. These methods optimize classifiers’ performance metrics, such as the True Positive rate (TPR) and false positive rate (FPR) [21, 22], and the most representative method is optimizing using the receiver operating characteristic curve (AUROC) [23]. These approaches find applications across disciplines, including econometrics [24], statistics [25], and machine learning [26]. While the AUROC-based method yields a globally optimized threshold tailored to the chosen model, it can lead to over-fitting or fail to adapt to other models.
Adaptive thresholding across groups:
Several studies advocate for group-adaptive decision thresholds to mitigate disparities [12, 27]. Menon and Williamson [28] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy. Similarly, Corbett-Davies et al. [29] demonstrate that group-specific thresholds can facilitate fairness-aware classification. Canetti et al. [30] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. Jang et al. [12] present threshold-adaptation methods to ensure fair classification, and Bakker et al. [27] explore threshold tuning to maintain stable classification performance. Building on these foundations, our framework FairOPT introduces a novel group-adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
Figure 3: Human-written short texts have a higher FPR compared to other text lengths, given the static threshold, showing the inadequacy of the static threshold to variations. Curves reflect the average probability of AI detectors in the training dataset.
3 Conceptual Framework and Methodology
This section outlines our method for adaptive thresholding, FairOPT, which iteratively adjusts classification thresholds for different subgroups to reduce discrepancy.
3.1 Objective
Consider a set of subgroups $G_{1},\ldots,G_{i}$ with attributes (e.g., length and style). Acknowledging that texts with varying characteristics show variations in probability distributions, our objective is to find optimal thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ for binary classification.
3.1.1 Preliminaries of AI Text Detection
The detection of AI-generated text is commonly formulated as a probabilistic classification for a binary classification, where the input space $\mathcal{X}$ consists of the sequence of tokens derived from a given text, and the label space $\mathcal{Y}=\{0,1\}$ denotes the two distinct classes of 00 for human-written and $1$ for AI-generated. To address this task, existing work primarily utilizes a pretrained neural network as a probabilistic classifier $M_{\theta}:\mathcal{X}\rightarrow[0,1]$, which assigns a probability $p$ to the input reflecting the likelihood of it being AI-generated [16, 13]. This probability is then converted into a predicted label by applying a threshold, typically set at $0.5$. Thus, the predicted label, AI-written, $\hat{Y}\in\{0,1\}$ is determined as follows:
$\hat{Y}=\begin{cases}1,&\text{if }M_{\theta}(x)\geq\text{threshold},\\ 0,&\text{otherwise.}\end{cases}$ (1)
3.1.2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix B.1.
3.1.3 Fairness metrics
To quantify disparity, we use Demographic Parity (DP) [31] and Balanced Error Rate (BER) [32]. DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equal across all groups. We define the DP metric:
$\Delta_{\text{DP}}$ $=\left|P(\hat{Y}=1\mid S=a)-P(\hat{Y}=1\mid S=b)\right|$ (2)
DP seeks to measure disparate impact, where one group receives disproportionate results [33], by addressing an overall equal rate of positive predictions, encompassing TP and TN, across groups. In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means no difference in overall TP and TN between groups.
In our subgroup-adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP. For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
$\mathrm{BER}=\tfrac{1}{2}\Bigl{(}\underbrace{\tfrac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}}_{\mathrm{FPR}}\;+\;\underbrace{\tfrac{\mathrm{FN}}{\mathrm{FN}+\mathrm{TP}}}_{\mathrm{FNR}}\Bigr{)}$ (3)
$\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$ (4)
Details about fairness metrics are noted in Appendix B.2.
Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section 3.2. Details about the tradeoff are in Appendix H.2.
3.2 Relaxed Fairness
While achieving $100\%$ fairness across groups (e.g., $DP=0$ across fifteen groups) is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria [34]. The $80\%$, relaxed fairness, rule has been suggested to offer a pragmatic approach (e.g., allowing $DP=0.2$ in achieving fairness) [33], which offers a more feasible fairness criterion and mitigates overfitting risks.
We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a given flexibility $\tau$, set to $0.8$:
$\frac{\min\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}{\max\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}\geq\tau,\quad\forall k\in\mathcal{K},$ (5)
where:
  • $\mathcal{S}$ denotes the set of attributes (e.g. length, writing style, etc.)
  • $\mathcal{M}$ represents the set of fairness metrics (e.g., demographic parity, equalized odds, etc.).
  • $c_{i}^{(k)}(\theta)$ computes the $k$-th metric in $\mathcal{K}$ for each group using classifier parameters $\theta$.
  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=1$ reflects 100% fairness, which is the most strict.
We limited the number of fairness metrics due to impossibility theorem [34] to make it computationally feasible, as it is impractical to satisfy all fairness criteria simultaneously.
Algorithm 1 FairOPT: Group-Specific Thresholding
Algorithm 1 FairOPT: Group-Specific Thresholding
1:  Input: $D$ dataset, $S_{n}$ features, predicted probabilities $p$, performance targets ($\alpha,\beta$), fairness gap limit $\delta_{\text{fair}}$, max iterations, step size $\eta$.
2:  Split $D$ into sub-datasets $D_{1},\ldots,D_{j}$ by $S_{n}$.
3:  for iteration = 1 to max iterations do
4:     for each subgroup $G_{i}$ do
5:        Convert $p$ in $G_{i}$ to binary predictions using $\theta(G_{i})$ and compute confusion matrix.
[…]
7:           return $\{\theta(G_{1}),\ldots,\theta(G_{j})\}$.
8:        else
9:           Adjust $\theta(G_{i})$
10:        end if
11:     end for
12:  end for
13:  Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
3.3 Group-Adaptive Threshold Optimization
Process
Given this setup, we introduce FairOPT  which is described in Algorithm 1. FairOPT uses a gradient-based subgroup threshold update to jointly optimize group-level fairness and classification performance. Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$. Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups). During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$. The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$. For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
$L_{i}(\theta)\;=\;-\,\mathrm{ACC}_{i}(\theta)\;+\;\kappa\,\bigl{[}\beta-\mathrm{F1}_{i}(\theta)\bigr{]}_{+},$ (6)
where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $[\cdot]_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
To optimize each $\theta(G_{i})$, we employ a finite-difference approximation of the gradient:
$\nabla L_{i}(\theta(G_{i}))\;\approx\;\frac{L_{i}(\theta(G_{i})+\delta)-L_{i}(\theta(G_{i})-\delta)}{2\,\delta}.$ (7)
The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the two loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ [35].
We then perform a gradient descent step for each subgroup:
$\theta(G_{i})\;\leftarrow\;\theta(G_{i})-\eta\,\nabla L_{i}(\theta(G_{i})),$ (8)
followed by clipping to $\bigl{[}a,b\bigr{]}\subseteq[0,1]$ to ensure a valid probability.
Finally, the model calculates the given fairness metric(s) $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the inter-group disparity:
$\Delta_{k}\;=\;\max_{i}\,M_{k}(G_{i})\;-\;\min_{i}\,M_{k}(G_{i}),$ (9)
and verifies whether $\Delta_{k}\leq\delta_{\mathrm{fair}}$ for each metric $M_{k}$. If any disparity exceeds $\delta_{\mathrm{fair}}$ (e.g., $0.2$), it is considered unfair. The fairness evaluation is not included in the main loss function because the discrepancy is evaluated in a binary manner: fair or unfair (e.g., pass or fail), similar to zero-one loss.
To determine termination, we monitor the largest threshold change among all groups,
$\Delta_{\theta}\;=\;\max_{G_{i}}\,\bigl{|}\theta_{\mathrm{new}}(G_{i})-\theta_{\mathrm{old}}(G_{i})\bigr{|}.$ (10)
[…]
Outcome
By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria. For full implementation details and hyperparameter specifications, refer to Appendix H.1.
Figure 5: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector results in cumulative probability distribution function (CDF; y-axis) based on the training dataset in the range of 0.1 to 0.9 (x-axis). The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$). Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions by stylistic features.
Figure 6: Threshold optimization based on the AUROC-based method under an $FPR\leq 10\%$ constraint (see Appendix C). The ROC curve (blue line) illustrates the trade-off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
4 Experiments
4.1 Dataset
Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [36], MAGE [37], and SemEval24 [38] with an even balance between human and AI text to avoid class imbalance. The generator includes GPT and Text-Davinci variants for their common usage, and the dataset spans a diverse range of writing domains, including social media (e.g., Reddit, ELI5), encyclopedic and academic sources (e.g., Wikipedia, arXiv), instructional and creative content (e.g., WikiHow, poetry, recipes), and reviews and summaries (e.g., Yelp, XSum), ensuring broad coverage of both formal and informal text styles. This experiment employed 19,117 instances for the training and 4,779 samples for the testing. We conducted the featured engineering noted in 4.3 and AI detections noted in 4.2 to extract features and get inferences from classifiers.
4.2 AI-Generated Text Classifiers
To classify AI-generated text, we employed multiple AI text detectors, DeskLib [39], E5LoRA [40], FastDetectGPT [41], RADAR [13], RoBERTa-base and -large [16], and SimpleAI’s detector [42] for open-source availability. Among closed-source models, GPT4.1-nano and GPT4o-mini [43, 44] have been used for LLM-based text detection [7]. These detectors were selected based on their common application and used in both the train and test datasets. The closed-source text detection is only conducted in the test dataset.
Figure 7: FairOPT decreases the Balanced Error Rate (BER) discrepancy of AI detectors compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$). The error bar shows the standard deviation from the mean.
4.3 Subgroup Attributes
We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics. These attributes are shown in Table 1. We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500). We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [45]. Details are noted in Appendix F. These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
Attribute Characteristics Labels
Text length Number of characters in text. Short, Medium, Long (k = 3)
Stylistic traits Inferred Big Five personality traits. Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism (k = 5)
Table 1: Subgroup attributes from the feature engineering.
Figure 8: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4. Our method incurs minimal loss of ACC, achieving $\mu=0.6344\pm 0.1518$ compared to Static ($\mu=0.6376\pm 0.1975$) and ROCFPR ($\mu=0.6249\pm 0.2057$) methods.
4.4 Discovery of Subgroup Bias in Text Detection
We assess detection performance across text-length subgroups under a static threshold $\theta=0.5$. As illustrated in Figure 3, short human-authored samples incur a substantially higher FPR than mid-length and long texts. This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
Next, we examine the combined effects of length and stylistic personality (Figure 5). By comparing the empirical cumulative distribution functions (CDFs) of AI-likelihood scores for short, conscientious versus long, neurotic texts, a two-sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix G.1.1 for full KS-test details). The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
4.5 Mitigating Subgroup Bias using FairOPT
Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups. We extracted the text length and personality features from every instance in the dataset, partitioning the corpus into 15 subgroups. For the dataset, multiple AI text classifiers were employed to generate predicted probabilities indicating the probability that a given text was AI-generated with $p\in[0,1]$.
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 2. We include all hyperparameters and training details in Appendix H.3. Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [16] rather than averaging the probabilities of multiple detectors. The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [46].
We also applied an AUROC-based thresholding method using the RoBERTa-large [16] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC-based optimization procedure [26]. The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score. Metrics are noted in Appendix B.2.
4.6 Tradeoff Analysis between Performance and Fairness
Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section 3.2 to less sacrificing ACC while still satisfying a demanded level of fairness.
The ROCFPR method, shown in Figure 6, demonstrates a notable ability to minimize the FPR in training. However, this reduction in FPR comes at a significant cost to the TPR. This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text, although the overall ACC is not overly sacrificed as shown in Figure 8.
4.7 Results with FairOPT for AI-Text Detection
Figure 7 demonstrates that group-adaptive thresholding from FairOPT consistently lowers the average discrepancy. Our method decreased the discrepancy to $0.4341$ ($\sigma=0.1927$), showing $12\%$ of reduction compared to the static-threshold baseline ($\mu=0.5512$, $\sigma=0.1758$) and outperforms the AUROC-tuned method ($\mu=0.5595$, $\sigma=0.1582$). Notably, E5LoRA shows the greatest decrease, with its BER discrepancy dropping by nearly half (from $0.63$ to $0.33$), and GPT4o-mini achieved a substantial reduction (from $0.60$ to $0.37$). The Desklib experiences a modest 0.003 reduction. Overall, the FairOPT decreased the discrepancy in all AI text classifiers. The detailed results with standard deviation by the dataset, detector, and thresholding methods are in Appendix LABEL:tab:overall_by_detector_and_dataset.
Figure 8 illustrates that FairOPT incurs only a 0.003 of sacrifice in overall ACC and reduction of the standard deviation by 0.04 while outperforming the AUROC-based method. These findings demonstrate that FairOPT substantially mitigates fairness disparities with negligible ACC tradeoff, offering a practical application. Notably, our method shows great effectiveness to commonly used AI detectors like DeskLib [39], FastDetectGPT [41], RADAR [13], and SimpleAI’s detector [42] by securing ACC over $75\%$ while reducing the discrepancy in all cases.
Moreover, the E5LoRA detector [40] exemplifies the ability of FairOPT to transcend the fairness–performance tradeoff. On the SemEval24 dataset, FairOPT achieves a 0.43 reduction in discrepancy while increasing ACC by 0.23. Similarly, on the MAGE, it delivers a 0.25 decrease in discrepancy coupled with a 0.21 ACC increase. These outcomes highlight the broad applicability of adaptive thresholding for improving both fairness and performance across diverse AI-text detection models, as well as other kinds of probabilistic classifiers.
5 Conclusion
This paper proposes FairOPT, a post-processing algorithm for probabilistic classifiers that reduces the discrepancy while preserving overall accuracy. Based on the observation that there is a statistically significant difference in AI detector outputs by text characteristics (e.g., length and personality), the need to address the variance has been raised because it can lead to notably higher error rates for specific groups. To address this problem, we proposed a new method for group-specific threshold optimization. Unlike universal thresholding methods, which presume evenly distributed probability regardless of the stylistic features of the text, our method effectively captures the variations of the probability distribution of AI text classifiers across different subgroups, leading to classification with a lower discrepancy.
While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio), including deepfake video and audio detectors. This method will be highly beneficial in post-processing, guiding the model toward decreased discrepancy, contingent upon the specific requirements and deployment context. We can also apply this method to all other probabilistic classifications when subgroup discrepancy matters (e.g., recidivism risk prediction or hiring decisions) by learning tailored decision thresholds that substantially mitigate bias with negligible loss in overall accuracy. Furthermore, while our primary focus has been improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training mechanism. Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.
Although our experiments focus on text-based detectors, the FairOPT framework is modality-agnostic and readily extends to other domains such as image, video, and audio deepfake detection, where probabilistic outputs require post-processing to ensure equitable performance. More generally, FairOPT applies to any probabilistic classification task with subgroup disparities (e.g., recidivism risk prediction, hiring decisions) by learning tailored decision thresholds that mitigate bias with minimal impact on overall accuracy.
6 Limitations
While FairOPT demonstrably reduces subgroup discrepancy on our English-text benchmarks, it relies on only two features (character-count bins and model-inferred Big Five traits). Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture. Moreover, the stylistic labels are derived from a downstream BERT-based personality predictor, which itself may introduce cultural bias. We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.
Impact Statement
This paper introduces novel theoretical frameworks and methodologies to enhance the robustness of AI-generated text detection systems. By extracting and using features for subgroups such as text length and stylistic traits, our approach addresses disparities in classification outcomes. This technology can play a crucial role in AI detection for identifying misinformation [3], safeguarding the integrity of publication and academic organizations [4, 6], and countering potential cybersecurity threats [5]. Since it is crucial for these AI-content detection methods to be robust and fair across many potential users, our method takes a major step in this direction by formulating the problem setting and developing a new algorithm.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes] We introduced a novel method that reduces the discrepancy of the probabilistic classifiers, especially in AI text detectors, and the main claim is noted both in the abstract and introduction.
  4. Justification: We noted the main claims in the abstract and the main body of the paper and provided experimental results as evidence to support the main claim.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes] Limitations are addressed at the end of the paper.
  9. Justification: Limitations about the AI text detection methods and questions around the inference to other kinds of probabilistic classifiers are addressed.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes] The paper provides the conceptual framework and the computational process
  14. Justification: We provides the confusion matrix, employed performance and fairness metrics, fairness criterion, multiple thresholding methods, and the optimization process based on the finite-difference approximation.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes] The paper includes the URL to the dataset and codes that are used for the main results.
  19. Justification: The paper provides the code and dataset. The experiment can be simulated with the provided dataset, and other experiments can also be conducted if they are based on the probabilistic classifiers and split into subgroups by adjusting the column names or adjusting hyperparameters.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes] We disclose all codes and dataset.
  24. Justification: You can access the data and code through the provided technical materials or via the URL.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes] We disclose the dataset balance, split, and characteristics of data. The employed hyperparameters are included in the appendix. The optimizer is developed in this research.
  29. Justification: The details about the dataset construction is noted in the dataset part, and appendix includes the hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes] We report the error bars, which implies the standard deviation of the results. Also, the statistical significance of the probability distribution is addressed using the K-S test in the main body of the paper.
  34. Justification: The reported error bar shows the variance of the result and we conducted an experiment to figure out the p-value of the statistical significance of the KS probability distribution.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No] This research is based on a local device (CPU) and does not use GPU
  39. Justification: The model is designed to be able to run in any environment without the constraint of the hardware.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes] This research complies with the Code of Ethics
  44. Justification: Since this research is not a human-subjects research, this project is exempt from the IRB investigation. This research does not include human-subject sensitive information, but they are augmented using feature engineering.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes] We have included potential impacts from this research in Section Impact Statement.
  49. Justification: This research generate sensitive features. But the extracted features do not infringe privacy because it is generating labels for each writing based on statistical inferences. This could lead a negative impacts. Regarding the positive impacts, the decreased discrepancy of the classifiers has been discussed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes] We used the clipping method to restrict the optimization of the probability in the range of 0 to 1
  54. Justification: Since the results outside of the range of 0 and 1 could lead to an error in the system, we used clipping in the model to prevent the exceed of the bound in the gradient descent process.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes] We used multiple AI text classifiers and they are all cited and addressed. Some codes may need the API keys from specific platforms like OpenAI
  59. Justification: The train and test datasets include multiple AI-generated text detectors. And the test dataset has two more detection results based on the open-source LLMs.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes] All resources used for the paper are provided in the URL
  64. Justification: We disclose all resources that used for the research.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] This research is not crowdsourcing research.
  69. Justification: The given multiple types of writings (e.g., essay, twitter, etc.) are constructed through the cited datasets, and the main novelty of the dataset is the application of multiple AI text detectors.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] This research is not human-subject research
  74. Justification: Some texts are written by human, but the sources are authorized and publicly accessible datasets.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes] This research used two language models for the AI text detection, which his not a conventional method despite there being research that shows its efficiency.
  79. Justification: This research used GPT4.1 and GPT4 like other AI text classifiers. This method is explored in other research as discussed in the main part of the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix B Performance and fairness metrics
B.1 Contingency table
The model can make not only true positive (TP) and true negative (TN) cases but also false positive (FP) and false negative (FN):
Classified as AI ($\hat{Y_{1}}$) Classified as non-AI ($\hat{Y_{0}}$)
AI-generated ($Y_{1}$) TP: Correct identification FN: Failed to detect AI and misidentified as human
Human-developed ($Y_{0}$) FP: Incorrectly marked human work as AI TN: Correctly identified by not classifying human work as AI
Table 2: Confusion matrix for the four possible outcomes when an AI detection system classifies either AI-generated or human-developed content.
B.2 Notation and Metrics
Given the table B.1 that denotes the outcomes in a binary classification task, we can mathematically express the performance and fairness metrics like below:
  • Accuracy (ACC).
    $\mathrm{ACC}\;=\;\frac{\mathrm{TP}+\mathrm{TN}}{\mathrm{TP}+\mathrm{TN}+\mathrm{FP}+\mathrm{FN}}.$
  • F1 Score. Given $\mathrm{Precision}$ and $\mathrm{Recall}$ below,
[…]
    F1 matters because it balances Precision and Recall into a single measure, making it especially useful when the class distribution is imbalanced or when both metrics are important.
  • Precision.
    $\mathrm{Precision}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}}$
    Precision measures the fraction of predicted positives that are actually correct, focusing on how often the model is right when it predicts a positive.
  • Recall.  (Recall is also referred to as TPR in some contexts.)
    $\mathrm{Recall}\;=\;\mathrm{TPR}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$
    Recall measures the fraction of actual positives that are correctly identified by the model, capturing how many real positives are detected.
  • Demographic Parity (DP). Suppose we have a protected attribute $A\in\{0,1\}$. Demographic Parity requires $P(\hat{Y}=1\mid A=0)=P(\hat{Y}=1\mid A=1)$, or
    $P(\hat{Y}=1\mid A=0)\approx P(\hat{Y}=1\mid A=1)$. A common metric is the disparity:
    $\Delta_{\mathrm{DP}}\;=\;\bigl{|}\,P(\hat{Y}=1\mid A=0)\;-\;P(\hat{Y}=1\mid A=1)\bigr{|}.$
  • Equalized Odds (EO). Equalized Odds requires equal true positive rates and false positive rates across groups. One way to measure the gap is:
    $\Delta_{\mathrm{EO}}\;=\;\bigl{|}P(\hat{Y}=1\mid Y=1,A=0)\;-\;P(\hat{Y}=1\mid Y=1,A=1)\bigr{|}\;+\;\ \$
    $\bigl{|}P(\hat{Y}=1\mid Y=0,A=0)\;-\;P(\hat{Y}=1\mid Y=0,A=1)\bigr{|}.$
    Under the strictest notion of fairness, one requires
[…]
    implying no disparity in true or false positive rates across groups.
    Relaxation in fairness allows
    $\Delta_{\mathrm{EO}}\geq 0,$
    for feasibility.
  • False Positive Rate (FPR).
    $\mathrm{FPR}\;=\;\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}$
  • Balanced Error Rate (BER).
    $\mathrm{BER}\;=\;\frac{1}{2}\left(\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}+\frac{\mathrm{FN}}{\mathrm{TP}+\mathrm{FN}}\right)\;=\;\frac{\mathrm{FPR}+\mathrm{FNR}}{2}\;=\;\frac{\mathrm{FPR}+(\mathrm{1-TPR})}{2}$
    Balanced Error Rate (BER) averages the error rates across both classes, providing a metric that is insensitive to class imbalance. By considering both the False Positive Rate and the False Negative Rate, BER ensures that the classifier performs equally well on both classes. The concept is especially well-introduced in [32]
  • BER Discrepancy.
    $\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$
    For each subgroup $G_{i}$ we can compute BERi to measure the discrepancy by comparing the maximum BER and minimum BER, and the lower discrepancy indicates more equitable classification across groups.
  • Specificity.
    $\mathrm{Specificity}\;=\;\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}}$
    Specificity measures the proportion of actual negative cases correctly identified by the model, reflecting its ability to detect true negatives. It is also known as true negative rate (TNR).
B.3 Impossibility theorem
The impossibility theorem typically refers to the result that no single classifier can satisfy all common fairness metrics (e.g., demographic parity, equalized odds, predictive parity) simultaneously, except in trivial cases. This arises because different fairness definitions can be mathematically incompatible when the base rates (prevalences) differ across groups. This research used one fairness metrics in the processing to address this issue.
Appendix C AUROC
C.1 Static threshold and AUROC-based approach
In this research, the AUROC-based approach is employed to optimize the classification threshold while balancing model performance and fairness constraints. This method evaluates the trade-off between the TPR and FPR, as visualized by the ROC curve shown in Figure 9. The optimal threshold is determined under a specified FPR constraint, $\text{FPR}\leq 0.1$, ensuring that the selected threshold aligns with a predefined ethical constraint.
Figure 9: Threshold optimization process using the AUROC-based approach with an FPR constraint of 0.1. The ROC curve (blue line) demonstrates the trade-off between TPR and FPR. The optimal threshold (green marker) is selected to maximize performance under the specified FPR constraint.
Appendix D Applied Thresholds
Figure 2 uses thresholds that is noted on Table 3.
Threshold Type Value
Static 0.5
ROCFPR 0.2866322
Adaptive Thresholds (n=15)
Group Threshold
long_agreeableness 0.988102
long_conscientiousness 0.983204
long_extroversion 0.991482
long_neuroticism 0.998293
long_openness 0.978291
medium_agreeableness 0.971192
medium_conscientiousness 0.980804
medium_extroversion 0.987482
medium_neuroticism 0.994291
medium_openness 0.969759
short_agreeableness 0.986492
short_conscientiousness 0.994242
short_extroversion 0.978118
short_neuroticism 0.994192
short_openness 0.970664
Table 3: Thresholds from three thresholding methods that used in the experiment.
Appendix E Detailed Results
We report a table showing the overall discrepancy, ACC, and F1 by dataset source: Table 4, and a table showing the detailed difference by detectors: Table LABEL:tab:overall_by_detector_and_dataset.
E.1 Overall discrepancy and performance
The table 4 summarizes the overall discrepancy and performance by each dataset and thresholding method. Detailed results that include the result from each detector is noted in the next section.
Source Strategy Discrepancy (mean $\pm$ std) Accuracy (mean $\pm$ std) F1 Score (mean $\pm$ std)
MAGE FairOPT $\mathbf{0.4495\pm 0.1656}$ $0.6339\pm 0.1512$ $0.3877\pm 0.3494$
ROCFPR $0.5551\pm 0.1664$ $0.6238\pm 0.2113$ $0.5326\pm 0.3170$
Static $0.5433\pm 0.1781$ $0.6388\pm 0.2015$ $0.5126\pm 0.3352$
RAID FairOPT $\mathbf{0.4836\pm 0.1781}$ $0.6280\pm 0.1573$ $0.3885\pm 0.3535$
ROCFPR $0.5644\pm 0.1263$ $0.6253\pm 0.2082$ $0.5363\pm 0.3185$
Static $0.5562\pm 0.1478$ $0.6364\pm 0.1994$ $0.5136\pm 0.3358$
SemEval FairOPT $\mathbf{0.3763\pm 0.2166}$ $0.6404\pm 0.1497$ $0.3896\pm 0.3484$
ROCFPR $0.5598\pm 0.1758$ $0.6258\pm 0.2010$ $0.5303\pm 0.3114$
Static $0.5548\pm 0.1969$ $0.6376\pm 0.1948$ $0.5081\pm 0.3288$
Table 4: Overall discrepancy and performance by data source and thresholding strategy (mean $\pm$ std)
E.2 Comparison by performance and fairness metrics by datasource, detector, and thresholding strategy
The table LABEL:tab:overall_by_detector_and_dataset shows the results by each detector by the datasource and thresholding method based on fairness and performance metrics.
Table 5: Discrepancy and performance (mean $\pm$ std) by data source, detector, and thresholding method
Dataset
Detector Thresholding $Discrepancy$ $ACC$ $F1score$
MAGE
GPT4.1-nano (2025) FairOPT $0.4106\pm 0.2350$ $0.4794\pm 0.0374$ $0.0114\pm 0.0121$
ROCFPR $0.4636\pm 0.2194$ $0.4274\pm 0.0380$ $0.0694\pm 0.0159$
Static $0.4647\pm 0.2143$ $0.4393\pm 0.0401$ $0.0507\pm 0.0154$
GPT4o-mini (2024) FairOPT $0.3820\pm 0.2421$ $0.4866\pm 0.0345$ $0.0046\pm 0.0078$
ROCFPR $0.6874\pm 0.1882$ $0.4453\pm 0.0288$ $0.2348\pm 0.0399$
Static $0.6517\pm 0.1661$ $0.4423\pm 0.0253$ $0.1947\pm 0.0304$
SimpleAI (2023) FairOPT $0.4536\pm 0.0871$ $0.7501\pm 0.0274$ $0.7006\pm 0.0363$
ROCFPR $0.5682\pm 0.0945$ $0.7573\pm 0.0130$ $0.7520\pm 0.0215$
Static $0.5527\pm 0.0968$ $0.7525\pm 0.0154$ $0.7360\pm 0.0298$
DeskLib (2024) FairOPT $0.4896\pm 0.0914$ $0.8900\pm 0.0170$ $0.8809\pm 0.0196$
ROCFPR $0.4854\pm 0.2388$ $0.8787\pm 0.0212$ $0.8916\pm 0.0224$
Static $0.4278\pm 0.2561$ $0.9074\pm 0.0206$ $0.9140\pm 0.0219$
E5LoRA (2024) FairOPT $0.3571\pm 0.2440$ $0.4895\pm 0.0336$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.5779\pm 0.1866$ $0.1996\pm 0.0302$ $0.0471\pm 0.0179$
Static $0.6026\pm 0.2439$ $0.2797\pm 0.0237$ $0.0065\pm 0.0090$
FastDetectGPT (2023) FairOPT $0.4929\pm 0.0189$ $0.8099\pm 0.0274$ $0.7774\pm 0.0243$
ROCFPR $0.4881\pm 0.0315$ $0.7776\pm 0.0302$ $0.8114\pm 0.0288$
Static $0.4911\pm 0.1295$ $0.8398\pm 0.0154$ $0.8520\pm 0.0151$
RoBERTa-base (2019) FairOPT $0.4893\pm 0.0537$ $0.6742\pm 0.0436$ $0.5381\pm 0.0655$
ROCFPR $0.5923\pm 0.0884$ $0.7322\pm 0.0311$ $0.6963\pm 0.0320$
Static $0.6158\pm 0.1606$ $0.7256\pm 0.0364$ $0.6722\pm 0.0423$
RoBERTa-large (2019) FairOPT $0.3940\pm 0.1992$ $0.5063\pm 0.0364$ $0.0656\pm 0.0372$
ROCFPR $0.5835\pm 0.0610$ $0.6264\pm 0.0289$ $0.5067\pm 0.0405$
Static $0.5414\pm 0.1038$ $0.6103\pm 0.0361$ $0.4315\pm 0.0522$
RADAR (2023) FairOPT $0.5760\pm 0.0755$ $0.6192\pm 0.0216$ $0.5106\pm 0.0388$
ROCFPR $0.5495\pm 0.2113$ $0.7693\pm 0.0186$ $0.7837\pm 0.0202$
Static $0.5418\pm 0.1280$ $0.7519\pm 0.0224$ $0.7562\pm 0.0242$
RAID
GPT4.1-nano (2025) FairOPT $0.5208\pm 0.2082$ $0.4738\pm 0.0087$ $0.0128\pm 0.0115$
ROCFPR $0.5820\pm 0.1623$ $0.4243\pm 0.0189$ $0.0795\pm 0.0295$
Static $0.5781\pm 0.1558$ $0.4306\pm 0.0183$ $0.0465\pm 0.0202$
GPT4o-mini (2024) FairOPT $0.4781\pm 0.2428$ $0.4710\pm 0.0180$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.5762\pm 0.1567$ $0.4557\pm 0.0266$ $0.2292\pm 0.0542$
Static $0.5816\pm 0.1733$ $0.4529\pm 0.0180$ $0.1875\pm 0.0321$
SimpleAI (2023) FairOPT $0.5868\pm 0.2080$ $0.7320\pm 0.0408$ $0.6835\pm 0.0548$
ROCFPR $0.5655\pm 0.1818$ $0.7452\pm 0.0386$ $0.7502\pm 0.0388$
Static $0.6245\pm 0.2095$ $0.7403\pm 0.0516$ $0.7357\pm 0.0541$
DeskLib (2024) FairOPT $0.3552\pm 0.1001$ $0.9023\pm 0.0289$ $0.8973\pm 0.0301$
ROCFPR $0.4578\pm 0.0522$ $0.8667\pm 0.0223$ $0.8843\pm 0.0194$
Static $0.4292\pm 0.0646$ $0.8988\pm 0.0212$ $0.9076\pm 0.0189$
E5LoRA (2024) FairOPT $0.4167\pm 0.2041$ $0.4808\pm 0.0153$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.6437\pm 0.1177$ $0.2059\pm 0.0384$ $0.0406\pm 0.0272$
Static $0.6535\pm 0.1981$ $0.2805\pm 0.0323$ $0.0078\pm 0.0097$
FastDetectGPT (2023) FairOPT $0.4405\pm 0.0560$ $0.8241\pm 0.0203$ $0.7996\pm 0.0230$
ROCFPR $0.5103\pm 0.0873$ $0.7850\pm 0.0345$ $0.8191\pm 0.0312$
Static $0.4609\pm 0.1409$ $0.8472\pm 0.0247$ $0.8601\pm 0.0228$
RoBERTa-base (2019) FairOPT $0.5361\pm 0.0452$ $0.6706\pm 0.0343$ $0.5483\pm 0.0681$
ROCFPR $0.6246\pm 0.1244$ $0.7348\pm 0.0355$ $0.7030\pm 0.0478$
Static $0.5978\pm 0.0984$ $0.7180\pm 0.0342$ $0.6682\pm 0.0499$
RoBERTa-large (2019) FairOPT $0.4282\pm 0.1927$ $0.4990\pm 0.0183$ $0.0698\pm 0.0285$
ROCFPR $0.6079\pm 0.0816$ $0.6413\pm 0.0218$ $0.5341\pm 0.0394$
Static $0.5570\pm 0.0658$ $0.6176\pm 0.0175$ $0.4579\pm 0.0319$
RADAR (2023) FairOPT $0.5900\pm 0.1850$ $0.5980\pm 0.0179$ $0.4853\pm 0.0412$
ROCFPR $0.5113\pm 0.0602$ $0.7683\pm 0.0263$ $0.7867\pm 0.0233$
Static $0.5232\pm 0.0644$ $0.7418\pm 0.0265$ $0.7510\pm 0.0257$
SemEval
GPT4.1-nano (2025) FairOPT $0.2663\pm 0.2792$ $0.4889\pm 0.0498$ $0.0110\pm 0.0116$
ROCFPR $0.5139\pm 0.2100$ $0.4459\pm 0.0354$ $0.0912\pm 0.0320$
Static $0.4889\pm 0.2156$ $0.4495\pm 0.0408$ $0.0503\pm 0.0210$
GPT4o-mini (2024) FairOPT $0.2763\pm 0.2923$ $0.4943\pm 0.0501$ $0.0029\pm 0.0076$
ROCFPR $0.5777\pm 0.2485$ $0.4638\pm 0.0350$ $0.2381\pm 0.0484$
Static $0.5617\pm 0.2843$ $0.4620\pm 0.0418$ $0.2008\pm 0.0364$
SimpleAI (2023) FairOPT $0.5655\pm 0.2123$ $0.7382\pm 0.0369$ $0.6842\pm 0.0260$
ROCFPR $0.6726\pm 0.1674$ $0.7430\pm 0.0313$ $0.7375\pm 0.0264$
Static $0.6519\pm 0.1904$ $0.7436\pm 0.0327$ $0.7283\pm 0.0231$
DeskLib (2024) FairOPT $0.3711\pm 0.0755$ $0.8942\pm 0.0349$ $0.8844\pm 0.0316$
ROCFPR $0.4487\pm 0.0727$ $0.8643\pm 0.0169$ $0.8776\pm 0.0241$
Static $0.3691\pm 0.0970$ $0.8954\pm 0.0139$ $0.9012\pm 0.0184$
E5LoRA (2024) FairOPT $0.2143\pm 0.2673$ $0.5015\pm 0.0523$ $00.0000\pm 0.0$ 0.0000
ROCFPR $0.6887\pm 0.2807$ $0.2098\pm 0.0405$ $0.0293\pm 0.0124$
Static $0.6428\pm 0.2790$ $0.2750\pm 0.0402$ $0.0065\pm 0.0061$
FastDetectGPT (2023) FairOPT $0.4818\pm 0.0312$ $0.8111\pm 0.0254$ $0.7701\pm 0.0298$
ROCFPR $0.4689\pm 0.0646$ $0.7490\pm 0.0284$ $0.7851\pm 0.0345$
Static $0.5232\pm 0.1360$ $0.8159\pm 0.0253$ $0.8268\pm 0.0276$
RoBERTa-base (2019) FairOPT $0.4922\pm 0.0426$ $0.6940\pm 0.0371$ $0.5676\pm 0.0472$
ROCFPR $0.5748\pm 0.1175$ $0.7519\pm 0.0241$ $0.7127\pm 0.0362$
Static $0.5543\pm 0.0866$ $0.7412\pm 0.0263$ $0.6847\pm 0.0435$
RoBERTa-large (2019) FairOPT $0.2956\pm 0.2093$ $0.5188\pm 0.0513$ $0.0718\pm 0.0086$
ROCFPR $0.5977\pm 0.1239$ $0.6479\pm 0.0218$ $0.5280\pm 0.0376$
Static $0.6073\pm 0.1346$ $0.6186\pm 0.0293$ $0.4333\pm 0.0302$
RADAR (2023) FairOPT $0.4241\pm 0.1244$ $0.6222\pm 0.0358$ $0.5140\pm 0.0407$
ROCFPR $0.4956\pm 0.0187$ $0.7561\pm 0.0310$ $0.7735\pm 0.0357$
Static $0.5939\pm 0.1847$ $0.7370\pm 0.0300$ $0.7407\pm 0.0361$
E.3 Results for Discrepancy, ACC, and F1
The bar charts illustrate the discrepancy, ACC, and F1 are incorporated in the main body of the paper, and this section encompasses a chart representing the F1 score, accompanied by three figures. The sequence of figures is arranged in the order of discrepancy, ACC, and F1.
Figure 10: FairOPT decreases the discrepancy, Balanced Error Rate (BER) discrepancy, of probabilistic classifiers compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.4341\pm 0.1927$, outperforming ROCFPR ($\mu=0.5595\pm 0.1582$) and Static ($\mu=0.5512\pm 0.1758$).
Figure 11: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness that is shown in Figure 4. Our method incurs only a negligible accuracy sacrifice, achieving $\mu=0.6344\pm 0.1518$ compared to Static $\mu=0.6376\pm 0.1975$ and ROCFPR $\mu=0.6249\pm 0.2057$.
Figure 12: Overall F1 have been sacrificed in the application of FairOPT. FairOPT achieves $\mu=0.389\pm 0.350$, compared to ROCFPR $\mu=0.533\pm 0.316$ and Static $\mu=0.511\pm 0.333$. Detectors with low ACC show a very low F1 score.
Appendix F Feature Engineering
In order to mitigate disparities within AI classifiers, a thorough feature engineering process was undertaken. Features were systematically extracted to characterize each text within the dataset. The primary focus of the experiment encompassed text length and anticipated personality traits. Although formality and sentiment were also extracted, these attributes are utilized to expedite processing, as the computational time increases exponentially with the inclusion of additional features. The subsequent table illustrates the features primarily extracted for the experiment, with an emphasis solely on the utilization of length and personality traits.
This research mainly used length and personality as a main feature in using FairOPT.
The text length attribute measures the number of characters in a text. OpenAI’s evaluations [26] revealed performance disparities between short (under 1,000 characters) and long texts when using their AI classifier. While their global threshold approach failed to address group-specific differences, it provided a valuable benchmark. Building on this and using exploratory data analysis (EDA), we refined the thresholds to 1,000 and 2,500 characters, categorizing texts as short (up to 1,000), mid-length (1,000–2,500), or long (over 2,500).
The personality attribute includes five traits: extroversion, neuroticism, agreeableness, conscientiousness, and openness. This attribute contributes to the enforcement of fairness metrics, such as demographic parity and equality of odds. For this feature, we determine the personality trait by selecting the one with the highest probability among the five using the [45] model, which is fine-tuned from BERT-base-uncased [47].
Appendix G Discrepancies by classifiers with additional features
G.1 Discrepancy of Probability in Subgroups
G.1.1 Discrepancy in RoBERTa-large Detector
The analysis revealed significant variability in the probability distributions between subgroup pairs, as depicted in Figure 13. Employing two-sample Kolmogorov-Smirnov (KS) statistic for two-sided tests, we assessed distributional discrepancies across various text groups characterized by distinct feature combinations, under the assumption of a non-parametric distribution. Significant differences were identified in certain subgroup comparisons, indicating notable discrepancies in their probability distributions. For example, comparing texts characterized by conscientiousness and short with those exhibiting neuroticism and long resulted in a KS statistic of 0.3081 ($p<0.01$). Similarly, evaluating conscientiousness and short against extroversion and medium yielded a KS statistic of 0.2616 ($p<0.01$). Conversely, smaller discrepancies were observed between agreeableness and short versus neuroticism and short (KS = 0.0909, $p<0.05$) and neuroticism and short versus openness and short (KS = 0.0842, $p<0.01$).
Figure 13: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoB33ERTa-large detector results in cumulative probability distribution function (CDF) based on the training dataset in the range of 0.1 to 0.9. The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$). Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions
Appendix H Operationalizing FairOPT
H.1 Algorithmic description
1:  Input: Labeled dataset $D=\{(x_{n},y_{n})\}$, predicted probabilities $\{p_{n}\}$, subgroup labels $\{g_{n}\}$ with possible groups $G_{1},\dots,G_{j}$, features $\{S_{1},\dots,S_{n}\}$, initial thresholds $\theta_{init}$, fairness metrics $\{M_{1},\ldots,M_{k}\}$ with acceptable disparity $\delta_{\mathrm{fair}}$, minimum performance thresholds: $\alpha$ (ACC) and $\beta$ (F1), penalty weight $\kappa$, learning rate $\eta$, tolerance $\epsilon_{\mathrm{tol}}$, maximum iterations, finite-difference step $\delta$, minimum probability for threshold $a$, maximum probability for threshold $b$
2:  Output: Subgroup thresholds $\{\theta^{*}(G_{1}),\dots,\theta^{*}(G_{j})\}$
3:  Initialization:
4:  split $D$ to ${(G_{1}),\dots,(G_{j})}$ by $S$
5:  while iteration $<$ maximum iterations do
6:     for each group $G_{j}$ do
7:        $\mathcal{I}_{i}\leftarrow\{\,n\mid g_{n}=G_{j}\}$  all samples in group $G_{j}$
8:        $\hat{y}_{n}\leftarrow[\,p_{n}\geq\theta(G_{j})\,],\;\forall n\in\mathcal{I}_{i}$
9:        Store contingency table for $G_{i}$ using $\hat{y}_{n}$ and ${y}_{n}$
10:        $\Omega\;\leftarrow\;\min_{i}\{\mathrm{ACC}(G_{j})\}\;\geq\;\alpha\;\wedge\;\min_{i}\{\mathrm{F1}(G_{j})\}\;\geq\;\beta$ boolean output
11:        for each fairness metric $M_{k}$ do
12:           Compute $\{M_{k}(G_{1}),\dots,M_{k}(G_{j})\}$
13:           $\Delta_{k}\leftarrow\max_{i}(M_{k}(G_{j}))-\min_{i}(M_{k}(G_{j}))$
14:           $\Psi\;\leftarrow\;\Bigl{[}\,\max_{k}\Delta_{k}\;\leq\;\delta_{\mathrm{fair}}\Bigr{]}$ boolean output
15:        end for
16:        for each group $G_{j}$ do
17:           $L_{i}(\theta)=-\mathrm{Acc}_{j}(\theta)+\kappa\,[\,\beta-\mathrm{F1}_{j}(\theta)\,]_{+}$
18:
$\nabla L_{i}(\theta(G_{j}))\;\approx\;\frac{L_{i}(\theta(G_{j})+\delta)\;-\;L_{i}(\theta(G_{j})-\delta)}{2\,\delta}.$
19:
$\theta(G_{j})\;\leftarrow\;\theta(G_{j})\;-\;\eta\;\nabla L_{i}(\theta(G_{j})).$
20:           Clip $\theta(G_{j})$ to stay within $[a,b]$
21:        end for
22:     end for
23:     $\Delta_{\theta}\;\leftarrow\;\max_{G_{j}}\bigl{|}\theta(G_{j})-\text{previousThresholds}[G_{j}]\bigr{|}$
24:     if ($\Delta_{\theta}<\epsilon_{\mathrm{tol}}\;\wedge\;\Omega\;\wedge\;\Psi$then
25:        break
26:     end if
[…]
28:     $\text{iteration}\leftarrow\text{iteration}+1$
29:  end while
30:  return $\{\theta(G_{1}),\dots,\theta(G_{j})\}$
Algorithm 2 FairOPT: Gradient-based adaptive threshold optimization with fairness for subgroups
H.2 Early stopping method for the tradeoff curve
The FairOPT algorithm was implemented using functions from the scipy.optimize library, aiming to determine probability thresholds that satisfy specific fairness constraints (acceptable_disparity = [1.00, 0.30, 0.25, 0.21, 0.20, 0.19, 0.15, 0.15]) while ensuring minimum performance thresholds for both performance criteria (ACC and F1). The optimization in the minimize function was performed using the L-BFGS-B method. The figure is based on the RoBERTa-base model, and it uses different hyperparameters to get eight results with different relaxation methods. Early stopping is facilitated by the default values of ftol and gtol in L-BFGS-B. Additionally, the learning rate was set to $10^{-3}$ to enable controlled and stable updates during optimization, and a penalty parameter was set to $10$ to strengthen the L2 penalty in the loss function. Furthermore, minimum performance thresholds for accuracy and F1-score were enforced based on the performance observed for each group when applying a static threshold of 0.5 and a ROC-based threshold. As stricter fairness constraints are imposed, these minimum thresholds are gradually relaxed.
H.3 Hyperparameter settings for the FairOPT
The application of the FairOPT is configured with a set of hyperparameters designed to ensure a balance between fairness and performance during the threshold optimization process. The learning rate is set to $10^{-3}$ to enable controlled and stable updates during optimization. The maximum number of iterations is specified as $10^{5}$ to provide sufficient time for the optimization algorithm to converge. An acceptable disparity of 0.2 is defined to regulate subgroup fairness, ensuring that performance differences across subgroups remain within an acceptable range through the relaxed fairness. Minimum thresholds for accuracy (0.25) and F1-score (0.25) are enforced to maintain baseline model performance and stability, particularly in cases of imbalanced datasets. A tolerance level of $1/2*10^{-4}$ is introduced to impose convergence criteria, ensuring the optimization process halts only when the changes in the objective function are negligible. Furthermore, a penalty parameter of 20 is applied to enforce performance constraints by penalizing significant performance disparities across subgroups. For fairness criteria, Demographic Parity (DP) was employed to test whether the generated thresholds meets the expected fairness. Also, this experiment used early stopping method for effective processing. The early stopping check through the change of the ACC, F1, and maximum DP discrepancy of each group, and terminates if all of them shows minimal change based on predefined patience.