### v5

# Minseok Jung1, Cynthia Fuertes Panizo2, Liam Dugan3, Yi R. (May) Fung1, Pin-Yu Chen4, Paul Pu Liang1

Abstract
The advancement of large language models (LLMs) has made it difficult to differentiate human-written text from AI-generated text. Several AI-text detectors have been developed in response, which typically utilize a fixed global threshold (e.g., $\theta=0.5$) to classify machine-generated text. However, one universal threshold could fail to account for distributional variations by subgroups. For example, when using a fixed threshold, detectors make more false positive errors on shorter human-written text, and more positive classifications of neurotic writing styles among long texts. These discrepancies can lead to misclassifications that disproportionately affect certain groups. We address this critical limitation by introducing FairOPT, an algorithm for group-specific threshold optimization for probabilistic AI-text detectors. We partitioned data into subgroups based on attributes (e.g., text length and writing style) and implemented FairOPT to learn decision thresholds for each group to reduce discrepancy. FairOPT showed notable discrepancy mitigation across nine detectors and three heterogeneous datasets, and the remarkable mitigation of the minimax problem by decreasing overall discrepancy 27.4% across five metrics while minimally sacrificing accuracy by 0.005%. Our framework paves the way for more robust classification in AI-generated content detection via post-processing. We release our data, code, and project information at URL.
I Introduction
The copious use of large language models (LLMs) facilitated widespread dissemination of AI-generated text [47, 20] and has raised concerns around the dissemination of misinformation [8], the degradation of publication standards [48], cybersecurity threats [49], and breaches of academic integrity [39]. In response to the anticipated risks, AI detection techniques have been proposed [47, 5]. Typically, these methods estimate a probability $p$ that the text is generated by AI by comparing the patterns (e.g., token probabilities) of the given text against AI-generated text, and use a universal threshold (e.g., $\theta>0.5$[17] for a binary classification.
Figure 1: We find that using one universal probability threshold (often $\theta>0.5$) to classify AI and human-generated text can fail to account for subgroup-specific distributional variations, such as text length or writing style, leading to higher errors for certain demographics. We propose FairOPT to learn multiple decision thresholds tailored to each subgroup, which reduces discrepancy in AI text detection as shown in Figure 2.
However, a single universal threshold for AI content classification neglects the heterogeneity of probability distributions by textual characteristics, such as length and writing style. To put in another way, one fixed threshold does not consider the variation of the probability from various stylistic features. For example, when using a fixed threshold, a writing style characterized by long neurotic writing could be more likely to be identified as AI-generated than one characterized by agreeableness, resulting in disproportionate misclassifications, affecting certain groups. Analogous disparities have been documented in other classification tasks, regarding race [1, 4] and gender [46, 25, 4].
To overcome these limitations, we introduce FairOPT, a group‐adaptive thresholding framework that learns distinct decision thresholds for subgroups rather than using methods with one universal threshold (see Figure 1). Our method balances performance and fairness metrics (e.g., demographic parity, or equality of odds) in learning thresholds. Experiment on nine AI text detectors on three benchmark datasets demonstrates that FairOPT substantially reduces discrepancy across subgroups while incurring only a marginal tradeoff in overall performance.
To summarize, our main contributions are: (a) identifying significant discrepancies in AI-generated text classifiers based on text length and stylistic characteristics, (b) developing FairOPT, a group-adaptive decision threshold that calibrates the decision boundaries to reduce group discrepancy, and (c) showing the effectiveness of the FairOPT, achieving overall 27.4% reduction across five metrics in discrepancy while showing negligible tradeoff in performance by dropping less than 0.1% of ACC.
II Literature review
AI text detectors
AI text detectors have been developed to tackle issues related to misinformation, integrity, and security. RADAR utilizes adversarial learning techniques [24], GLTR employs statistical metrics, including entropy, probability distribution over the model’s vocabulary for each next token (i.e. the softmax scores across all candidate tokens), and records the rank position of the actual chosen token within that sorted probability to discern [18], DetectGPT examines the curvature of log probability distributions of the referred LLM, concentrating on local likelihood curvature to detect AI-generated text [35], and RoBERTa-based detectors fine-tune pretrained language models [42]. We employed probabilistic classifiers for an AI‐likelihood score that use a threshold for classification rather than architectures that generate a direct discrimination.
Static thresholding for classification
The canonical thresholding in probabilistic classification is using a threshold of $0.5$ to map the probabilities to labels [17]. This default rule is commonly used in many fields like photogrammetry [41], ecology [32, 21], and machine learning [31]. Nevertheless, as highlighted by [17], the dependence solely on $\theta=0.5$ as a default threshold, merely due to its conventional acceptance, could be unreliable because of variations in data.
Universal thresholding with optimization
Threshold optimization techniques have been formulated to learn an optimal threshold across the dataset. These methods optimize classifiers’ performance metrics, such as the True Positive rate (TPR) and false positive rate (FPR) [27, 29], and the most representative method is optimizing using the receiver operating characteristic curve (AUROC) [6]. These approaches find applications across disciplines, including econometrics [43], statistics [14], and machine learning [36]. While the AUROC-based method yields a globally optimized threshold tailored to the chosen model, it can lead to over-fitting or fail to adapt to other models.
Adaptive thresholding across groups
Several studies advocate for group‐adaptive decision thresholds to mitigate disparities [25, 2]. [33] propose instance-dependent thresholds for inferred probabilities, resulting in reduced discrepancy. Similarly, [9] demonstrate that group-specific thresholds can facilitate fairness-aware classification. [7] investigate thresholds differentiated by race and proved that adaptive thresholds reduce performance disparity. [25] present threshold-adaptation methods to ensure fair classification, and [2] explore threshold tuning to maintain stable classification performance. Building on these foundations, our framework FairOPT introduces a novel group‐adaptive thresholding algorithm that jointly optimizes performance and fairness objectives.
III Conceptual Framework and Methodology
Figure 2: FairOPT decreases discrepancy of AI detectors compared to other thresholding methods. FairOPT reduces classification discrepancy to $\mu=0.613$, outperforming ROCFPR ($\mu=0.920$) and Static ($\mu=0.887$). The error bar shows the standard deviation from the mean. Each bar shows average of five types of discrepancies that has been introduced in III-A4.
Figure 3: Human-written short texts have a higher FPR compared to other text lengths, given the static threshold, showing the inadequacy of the static threshold to variations. Curves reflect the average probability of AI detectors in the training dataset.
TABLE I: FairOPTsignificantly reduces subgroup discrepancies across all evaluation metrics while keeping accuracy on par with baselines. This table compares thresholding methods: FairOPT, ROCFPR, and Static on ACC and five fairness metrics that measure discrepancies.
Metric FairOPT ROCFPR Static
ACC 0.6341 0.6249 0.6376
MaxBER Discrepancy 0.5852 0.7592 0.7507
Demographic Parity Difference 0.6772 0.9826 0.9730
Equalized Odds Difference 0.6379 0.9198 0.8501
FPR Discrepancy 0.4313 0.9715 0.9471
Predictive Parity Difference 0.7333 0.9684 0.9130
TABLE II: FairOPT achieves the lowest subgroup discrepancies among three thresholding methods while maintaining accuracy. Comparison of thresholding methods has been made across three datasets: MAGE (MG), RAID (RD), and SemEval (SE) and their overall averages (All). MaxBER = BER Discrepancy, MaxDP = Demographic Parity Discrepancy, MaxEO = Equalized Odds Discrepancy, MaxFPR = False Positive Rate Discrepancy, MaxPP = Predictive Parity Discrepancy. The discrepancy based on the MiniMax problem has been noted in III-A4.
Method MaxBER Disc. MaxDP Disc. MaxEO Disc.
All MG RD SE All MG RD SE All MG RD SE
Static 0.751 0.730 0.771 0.754 0.973 0.972 0.958 0.987 0.850 0.859 0.838 0.852
ROCFPR 0.759 0.745 0.783 0.753 0.983 0.978 0.977 0.992 0.920 0.949 0.901 0.906
FairOPT 0.585 0.582 0.582 0.591 0.677 0.671 0.659 0.700 0.638 0.635 0.629 0.649
Method MaxFPR Disc. MaxPP Disc. Accuracy
All MG RD SE All MG RD SE All MG RD SE
Static 0.947 0.939 0.974 0.932 0.913 0.902 0.920 0.918 0.638 0.639 0.636 0.638
ROCFPR 0.972 0.971 0.985 0.961 0.968 0.967 0.982 0.958 0.625 0.624 0.625 0.626
FairOPT 0.431 0.444 0.416 0.432 0.733 0.746 0.713 0.738 0.634 0.633 0.628 0.640
This section outlines our method for adaptive thresholding, FairOPT, which iteratively adjusts classification thresholds for different subgroups to reduce discrepancy.
III-A Objective
Consider a set of subgroups $G_{1},\ldots,G_{i}$ with attributes (e.g., length and style). Acknowledging that texts with varying characteristics show variations in probability distributions, our objective is to find optimal thresholds $\{\theta(G_{1}),\ldots,\theta(G_{i})\}$ for binary classification.
III-A1 Preliminaries of AI Text Detection
[…]
$\hat{Y}=\begin{cases}1,&\text{if }M_{\theta}(x)\geq\text{threshold},\\ 0,&\text{otherwise.}\end{cases}$ (1)
III-A2 Performance metrics
After inferring $\hat{Y}$, we can compare $Y$ and $\hat{Y}$. Details about the confusion matrix and performance metrics are in Appendix -A.
III-A3 Minimax problem
We quantify subgroup disparity using five complementary metrics—Demographic Parity (DP) [26], Balanced Error Rate (BER) [16], Equalized Odds, FPR, and Predictive Parity. For each metric $k$, let $m_{k,g}$ denote the value on subgroup $g$; we define the inter-group discrepancy $\Delta_{k}\!=\!\max_{g}m_{k,g}\;-\;\min_{g}m_{k,g}$ (lower is better). We then adopt a minimax objective,
$\min_{\theta}\;\max_{k}\;\Delta_{k}(\theta),$
because fairness guarantees should hold in the worst subgroup rather than only on average. Controlling $\max_{k}\Delta_{k}$ addresses tail risk and improves robustness.
III-A4 Fairness metrics
DP is a fairness criterion measuring whether a classifier’s predictions $\hat{Y}$ are statistically independent across attributes $S$: $\hat{Y}\perp\!\!\!\perp S$. This condition mandates that the probability of assigning a positive outcome is equal across all groups. We define the DP metric:
$\Delta_{\text{DP}}$ $=\left|P(\hat{Y}=1\mid S=a)-P(\hat{Y}=1\mid S=b)\right|$ (2)
DP seeks to measure disparate impact, where one group receives disproportionate results [15], by addressing an overall equal rate of positive predictions, encompassing TP and TN, across groups. In applying DP, the perfect fairness is $\Delta_{\text{DP}}=0$, which means no difference in overall TP and TN between groups.
In our subgroup‐adaptive evaluation, we use BER to equally measure the type I and II errors, which are FP and FN, while including positive predictions from DP. For each subgroup $G_{i}$ we compute its BERi to measure the discrepancy, and the lower discrepancy indicates more equitable classification across groups.
$\mathrm{BER}=\tfrac{1}{2}\Bigl(\underbrace{\tfrac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TN}}}_{\mathrm{FPR}}\;+\;\underbrace{\tfrac{\mathrm{FN}}{\mathrm{FN}+\mathrm{TP}}}_{\mathrm{FNR}}\Bigr)$ (3)
$\text{Discrepancy}=\max\mathrm{BER}_{i}\;-\;\min\mathrm{BER}_{i}$ (4)
Discrepancy has been measured based on minimax problem to minimize the maximum subgroup discrepancy (e.g., BER gap, FPR gap), thereby reducing the worst-case disparity.
Equalized Odds (EO) requires that a classifier’s predictions are conditionally independent of sensitive attributes $S$ given the true label $Y$ [22]. This means that both the TPR and FPR should be equal across groups. In practice, EO measures whether individuals from different subgroups are treated similarly in terms of both favorable outcomes (correctly identifying positives) and unfavorable outcomes (incorrectly labeling negatives as positives). The discrepancy is defined as the maximum gap across subgroups in either TPR or FPR:
$\Delta_{\text{EO}}$ $=\max\Bigl(\max_{g}\text{TPR}-\min_{g}\text{TPR},\;\max_{g}\text{FPR}-\min_{g}\text{FPR}\Bigr).$ (5)
Perfect fairness under EO is achieved when $\Delta_{\text{EO}}=0$, meaning that subgroup differences in both detection sensitivity and error rates are eliminated.
False Positive Rate (FPR) Discrepancy focuses specifically on subgroup differences in the rate of false alarms, i.e., when the classifier predicts $\hat{Y}=1$ despite the true label being $Y=0$. The discrepancy is formally defined as the difference between the maximum and minimum false positive rates across groups:
$\Delta_{\text{FPR}}$ $=\max_{g}\text{FPR}_{g}-\min_{g}\text{FPR}_{g}.$ (6)
This metric captures whether some groups are disproportionately subjected to incorrect positive predictions, which is critical in high-stakes applications such as fraud detection.
Predictive Parity (PP) Discrepancy measures whether a positive prediction carries the same meaning for different groups, i.e., whether $P(Y=1\mid\hat{Y}=1,S=g)$ is group-independent. The discrepancy is defined as:
$\Delta_{\text{PP}}$ $=\max_{g}\text{PP}_{g}-\min_{g}\text{PP}_{g},$ (7)
where $\text{PP}_{g}=\frac{\text{TP}_{g}}{\text{TP}_{g}+\text{FP}_{g}}$. PP evaluates whether the precision of a classifier, defined as the proportion of true positives among all predicted positives, is consistent across subgroups.
Addressing the mini–max problem in fairness is important because it explicitly targets the worst-case disparity across subgroups rather than optimizing only for average performance. Details about fairness metrics are noted in Appendix -B.
Figure 4: The FairOPT algorithm enables us to obtain a balance in the tradeoff between performance and fairness, through a relaxed fairness criterion discussed in Section III-B. Details about the tradeoff are in Appendix -H.
III-B Relaxed Fairness
While achieving $100\%$ fairness across groups (e.g., $DP=0$ across fifteen groups) is theoretically ideal, it is often impractical due to computational constraints and the difficulty of ensuring model convergence under strict fairness criteria [13]. The $80\%$, relaxed fairness, rule has been suggested to offer a pragmatic approach (e.g., allowing $DP=0.2$ in achieving fairness) [15], which offers a more feasible fairness criterion and mitigates overfitting risks.
We formalize a relaxed fairness criterion requiring that the ratio of minimum to maximum value of each fairness metric across all subgroups does not fall below a given flexibility $\tau$, set to $0.8$:
$\frac{\min\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}{\max\limits_{i\in\mathcal{S}}\,c_{i}^{(k)}(\theta)}\geq\tau,\quad\forall k\in\mathcal{K},$ (8)
where:
  • $\mathcal{S}$ denotes the set of attributes (e.g. length, writing style, etc.)
  • $\mathcal{M}$ represents the set of fairness metrics (e.g., demographic parity, equalized odds, etc.).
  • $c_{i}^{(k)}(\theta)$ computes the $k$-th metric in $\mathcal{K}$ for each group using classifier parameters $\theta$.
  • $\tau\in[0,1]$ is the fairness criteria, with $\tau=1$ reflects 100% fairness, which is the most strict.
We limited the number of fairness metrics due to impossibility theorem [13] to make it computationally feasible, as it is impractical to satisfy all fairness criteria simultaneously.
Algorithm 1 FairOPT: Group-Specific Thresholding
1:Input: $D$ dataset, $S_{n}$ features, predicted probabilities $p$, performance targets ($\alpha,\beta$), fairness gap limit $\delta_{\text{fair}}$, max iterations, step size $\eta$.
2: Split $D$ into sub-datasets $D_{1},\ldots,D_{j}$ by $S_{n}$.
3:for iteration = 1 to max iterations do
4:  for each subgroup $G_{i}$ do
5:   Convert $p$ in $G_{i}$ to binary predictions using $\theta(G_{i})$ and compute confusion matrix.
6:   if ($\mathrm{ACC}_{i}\geq\alpha$ and $\mathrm{F1}_{i}\geq\beta$ for all ii) and (fairness gap $\leq\delta_{\text{fair}}$) then
7:    return $\{\theta(G_{1}),\ldots,\theta(G_{j})\}$.
8:   else
9:    Adjust $\theta(G_{i})$
10:   end if
11:  end for
12:end for
13:Output: $\{\theta(G_{1}),\dots,\theta(G_{j})\}$.
III-C Group-Adaptive Threshold Optimization
Process
Given this setup, we introduce FairOPT  which is described in Algorithm 1. FairOPT uses a gradient-based subgroup threshold update to jointly optimize group-level fairness and classification performance. Formally, we partition the dataset $D$ into subgroups $\{G_{1},\dots,G_{i}\}$ based on features $\{S_{1},\dots,S_{n}\}$. Each subgroup $G_{i}$ is assigned an initial threshold $\theta_{\mathrm{init}}$ (e.g., 0.5 to all groups). During each iteration, we compute predictions for each sample $x_{n}\in G_{i}$ by binarizing the predicted probability $p_{n}$ using the current threshold $\theta(G_{i})$. The model derives a contingency matrix based on the predictions $\hat{y}_{n}$ and actual values $y_{n}$. For each subgroup $G_{i}$, we define a loss function that penalizes both low accuracy and insufficient F1 score:
$L_{i}(\theta)\;=\;-\,\mathrm{ACC}_{i}(\theta)\;+\;\kappa\,\bigl[\beta-\mathrm{F1}_{i}(\theta)\bigr]_{+},$ (9)
where $\beta$ is a minimum F1, $\kappa>0$ is a penalty weight, $\alpha$ is given for the minimum ACC, and $\cdot$$_{+}$ denotes the positive part that only positive values are retained, turning negative numbers to zero. These minimum standards are essential for ensuring stable convergence and preventing degenerate outcomes across all subgroups.
To optimize each $\theta(G_{i})$, we employ a finite-difference approximation of the gradient:
$\nabla L_{i}(\theta(G_{i}))\;\approx\;\frac{L_{i}(\theta(G_{i})+\delta)-L_{i}(\theta(G_{i})-\delta)}{2\,\delta}.$ (10)
The finite-difference approximation estimates the gradient of each subgroup’s threshold-based loss function by computing the difference between the two loss values at $\theta(G_{i})+\delta$ and $\theta(G_{i})-\delta$, then dividing by $2\delta$ [30].
We then perform a gradient descent step for each subgroup:
$\theta(G_{i})\;\leftarrow\;\theta(G_{i})-\eta\,\nabla L_{i}(\theta(G_{i})),$ (11)
followed by clipping to $\bigl[a,b\bigr]\subseteq[0,1]$ to ensure a valid probability.
Finally, the model calculates the given fairness metric(s) $\{M_{1},\ldots,M_{k}\}$ across all subgroups by measuring the inter-group disparity:
$\Delta_{k}\;=\;\max_{i}\,M_{k}(G_{i})\;-\;\min_{i}\,M_{k}(G_{i}),$ (12)
and verifies whether $\Delta_{k}\leq\delta_{\mathrm{fair}}$ for each metric $M_{k}$. If any disparity exceeds $\delta_{\mathrm{fair}}$ (e.g., $0.2$), it is considered unfair. The fairness evaluation is not included in the main loss function because the discrepancy is evaluated in a binary manner: fair or unfair (e.g., pass or fail), similar to zero-one loss.
To determine termination, we monitor the largest threshold change among all groups,
$\Delta_{\theta}\;=\;\max_{G_{i}}\,\bigl|\theta_{\mathrm{new}}(G_{i})-\theta_{\mathrm{old}}(G_{i})\bigr|.$ (13)
If $\Delta_{\theta}$ falls below a tolerance $\epsilon_{\mathrm{tol}}$ and all groups meet the performance criterion ($\mathrm{ACC}_{i}\geq\alpha$, $\mathrm{F1}_{i}\geq\beta$) and fairness disparities ($\Delta_{k}\leq\delta_{\mathrm{fair}}$) hold, the algorithm terminates.
Outcome
By iterating this process, the Algorithm 1 converges to an optimized set of subgroup thresholds $\{\theta(G_{1}),\dots,\theta(G_{i})\}$ that reduces the disparity while satisfying the demanded performance criteria. For full implementation details and hyperparameter specifications, refer to Appendix -G.
Figure 5: Top two highest and lowest distributional differences using Kolmogorov-Smirnoff (KS) test on RoBERTa-large detector results in cumulative probability distribution function (CDF; y-axis) based on the training dataset in the range of 0.1 to 0.9 (x-axis). The biggest discrepancy is observed with a KS statistic of $0.3081$ ($p<0.01$), while the smallest is $0.0842$ ($p<0.01$). Statistically significant divergences between specific groups call attention to the need for thresholding methods that are accountable to the varying distributions by stylistic features.
Figure 6: Threshold optimization based on the AUROC‐based method under an $FPR\leq 10\%$ constraint (see Appendix AUROC). The ROC curve (blue line) illustrates the trade‐off between TPR and FPR, and the optimal threshold (green marker) maximizes TPR under the specified FPR constraint.
IV Experiment
IV-A Dataset
Our experiments aim to assess the performance of FairOPT in comparison to the static and the AUROC-based threshold optimization methods The dataset is constructed with the RAID [12], MAGE [28], and SemEval24 [45] with an even balance between human and AI text to avoid class imbalance. The generator includes GPT and Text-Davinci variants for their common usage, and the dataset spans a diverse range of writing domains, including social media (e.g., Reddit, ELI5), encyclopedic and academic sources (e.g., Wikipedia, arXiv), instructional and creative content (e.g., WikiHow, poetry, recipes), and reviews and summaries (e.g., Yelp, XSum), ensuring broad coverage of both formal and informal text styles. This experiment employed 19,117 instances for the training and 4,779 samples for the testing. We conducted the featured engineering noted in IV-C and AI detections noted in IV-B to extract features and get inferences from classifiers.
IV-B AI-Generated Text Classifiers
To classify AI-generated text, we employed multiple AI text detectors, DeskLib [10], E5LoRA [50], FastDetectGPT [3], RADAR [24], RoBERTa-base and -large [42], and SimpleAI’s detector [19] for open-source availability. Among closed-source models, GPT4.1-nano and GPT4o-mini [38, 37] have been used for LLM-based text detection [5]. These detectors were selected based on their common application and used in both the train and test datasets. They are based on probabilistic inference framework that needs a threshold for the final classification. The closed-source text detection is only conducted in the test dataset.
IV-C Subgroup Attributes
We extracted attributes from each text for sub-grouping to capture heterogeneity in text characteristics. These attributes are shown in Table III. We categorized text length into three classes based on the number of characters as short (up to 1,000), mid-length (1,000 to 2,500), or long (over 2,500). We also labeled text with stylistic personality based on five traits by selecting the one with the highest probability among the five using a BERT-based personality prediction model [34]. The third feature is formality that measures the formal expression of the sentence based on Heylighen & Dewaele’s formality score [23] and two categories (formal and informal) are labeled. The fourth feature is a binary sentiment label that has been inferred using DistilBERT-based SST-2 classifier[40]. Details are noted in Appendix Feature Engineering. These features were chosen because text length captures structural variation and stylistic personality reflects writing tendencies, both of which critically influence detector stability.
TABLE III: Subgroup attributes from the feature engineering.
Attribute Characteristics Labels
Text length Number of characters in text. Short, Medium, Long (k=3)
Stylistic traits Inferred Big Five personality traits. Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism (k=5)
Formality Formality of the writing Formal, informal (k=2)
Sentiment Binary sentiment Positive, negative (k=2)
Figure 7: FairOPT shows a negligible sacrifice of ACC by effectively handling the performance-fairness tradeoff using a relaxed fairness rule that is shown in Figure 4. Our method incurs minimal loss of ACC, achieving $\mu=0.6341$ compared to Static ($\mu=0.6249$) and ROCFPR ($\mu=0.6376$) methods.
IV-D Discovery of Subgroup Bias in Text Detection
We assess detection performance across text‐length subgroups under a static threshold $\theta=0.5$. As illustrated in Appendix Detailed Results, short human‐authored samples incur a substantially higher FPR than mid‐length and long texts. This length-dependent disparity underscores the necessity of text length adaptive thresholding to mitigate the discrepancy.
Next, we examine the combined effects of length and stylistic personality (Figure 5). By comparing the empirical cumulative distribution functions (CDFs) of AI‐likelihood scores for short, conscientious versus long, neurotic texts, a two‐sample Kolmogorov–Smirnov (KS) test reveals a statistically significant divergence in their score distributions ($p<0.05$), indicating that static thresholding disproportionately misclassifies certain stylistic and length combinations (see Appendix -F1 for full KS‐test details). The observed disparity in the combination of text length and stylistic personality further highlights the need for adaptive thresholding to guarantee equitable detection performance across heterogeneous text cohorts.
IV-E Mitigating Subgroup Bias using FairOPT
Given these findings, we now describe the experimental setup for applying FairOPT for AI-text detection with reduced discrepancy across subgroups. We extracted the textual features from every instance in the dataset, partitioning the corpus into 60 subgroups. Across each dataset, multiple AI text classifiers were used to infer the probability $p\in[0,1]$ that a given text was AI-generated.
We used FairOPT to get adaptive thresholds $\{\theta(G_{1}),\dots,\theta(G_{j})\}$ for each subgroup and applied them as noted in Figure 9. We include all hyperparameters and training details in Appendix -I. Threshold optimization was performed on the training dataset based on the results from RoBERTa-large [42] rather than averaging the probabilities of multiple detectors. The relaxed fairness criterion is integrated to encourage earlier convergence and minimize the sacrifice of overall performance while pursuing group fairness [44].
We also applied an AUROC‐based thresholding method using the RoBERTa‐large [42] to select the decision threshold that maximizes the TPR under the constraint $FPR<0.1$; see Figure 6 for the AUROC‐based optimization procedure [36]. The optimized thresholds were then applied to the test dataset and evaluated based on discrepancy, ACC, and F1 score. Metrics are noted in Appendix -B.
IV-F Tradeoff Analysis between Performance and Fairness
Figure 4 shows the performance-fairness tradeoff by our proposed FairOPT algorithm compared to the Static and ROCFPR baselines. The relaxed fairness criterion from Section III-B to less sacrificing ACC while still satisfying a demanded level of fairness.
The ROCFPR method, shown in Figure 6, demonstrates a notable ability to minimize the FPR in training. However, this reduction in FPR comes at a significant cost to the TPR. This trade-off is particularly problematic as it diminishes the detector’s reliability in correctly identifying AI-generated text, although the overall ACC is not overly sacrificed as shown in Figure 7.
IV-G Results with FairOPT for AI-Text Detection
Table I shows that group-adaptive thresholding with FairOPT consistently reduces subgroup discrepancy across detectors. Overall, our method decreased the average discrepancy around $27.4\%$, substantially reducing subgroup gaps compared to static and ROCFPR baselines. On average, FairOPT achieves a BER discrepancy of $0.5852$, representing a $22\%$ reduction relative to the static threshold baseline ($0.7507$) and a $23\%$ reduction compared to ROCFPR ($0.7592$). Similar improvements are observed across other metrics: demographic parity decreases from $0.9730$ to $0.6772$ ($30\%$), equalized odds from $0.8501$ to $0.6379$ ($25\%$), false-positive rate discrepancy from $0.9471$ to $0.4313$ ($51\%$), and predictive parity from $0.9130$ to $0.7333$ ($18\%$). Overall, FairOPT lowers the discrepancy. Full results, including means and standard deviations by dataset, detector, and thresholding method, are provided in Appendix LABEL:tab:overall_by_detector_and_dataset.
Regarding the consistency of FairOPT across datasets, it successfully reduces discrepancy across all fairness metrics compared to both Static and ROCFPR, as shown in Table II. On average, the MaxBER discrepancy decreases by $22\%$, dropping from 0.751 (Static) and 0.759 (ROCFPR) to 0.585. Similarly, the MaxDP discrepancy decreases by $30\%$, from 0.973 (Static) and 0.983 (ROCFPR) to 0.677, while the MaxEO discrepancy decreases by $25\%$, from 0.850 (Static) and 0.920 (ROCFPR) to 0.638. The strongest improvement is observed for MaxFPR, where FairOPT achieves a $55\%$ reduction, falling from 0.947 (Static) and 0.972 (ROCFPR) to 0.431. Predictive Parity discrepancy also improves by $20\%$, decreasing from 0.913 (Static) and 0.968 (ROCFPR) to 0.733. Importantly, these fairness gains come with negligible accuracy cost: overall accuracy remains stable at 0.634 for FairOPT, essentially unchanged from 0.638 under Static and higher than 0.625 under ROCFPR.
Across detectors, FairOPT consistently reduces the average discrepancy as shown in 2 while making a negligible sacrifice of ACC 7. Across detectors, FairOPT lowers the mean discrepancy to 0.613 versus 0.887 (static) and 0.920 (ROCFPR). E5LoRA shows the largest gain, dropping to 0.100 from 0.672 (static) and 0.878 (ROCFPR). Among BERT-based classifiers, RoBERTa-base falls to 0.782 from 0.927 (static) and 0.932 (ROCFPR), while RoBERTa-large drops to 0.490 from 0.890 (static) and 0.940 (ROCFPR), both indicating substantial reductions. GPT-based detectors also improve meaningfully: GPT4o-mini to 0.253 and GPT4.1-nano to 0.466, compared to 0.9 under both static and AUROC. Other models show smaller but positive gains—DeskLib 12%, FastDetectGPT 10%, and SimpleAI 5%. The lone exception is RADAR, which increases by 2% under FairOPT; here, AUROC is approximately 1% lower than static. Detailed results with standard deviations are included in Appendix LABEL:tab:overall_by_detector_and_dataset.
Figure 7 shows that the fairness improvements achieved by FairOPT come with only negligible predictive cost. Averaged across all datasets, FairOPT achieves an accuracy of $0.6341$, compared to $0.6376$ for Static and $0.6249$ for ROCFPR. The difference compared to static is less than $0.1\%$. These findings confirm that FairOPT substantially mitigates fairness disparities with negligible accuracy tradeoff, making it a practical approach for fair AI-text detection.
IV-H Threshold Stability
The stability of FairOPT were systematically evaluated across three highly diverse and representative benchmark datasets: MAGE, RAID, and SemEval. These datasets were intentionally selected to capture a broad spectrum of domains, genres, and writing styles, including academic and encyclopedic texts, social media posts, creative writing, and reviews. This diversity ensures that our subgroup definitions, based on features such as length, personality, formality, and sentiment, are well supported and meaningful within each dataset. This provides a rigorous testbed for threshold optimization.
V Conclusion
This paper proposes FairOPT, a post-processing algorithm for probabilistic classifiers that reduces the discrepancy while preserving overall accuracy. Based on the observation that there is a statistically significant difference in AI detector outputs by text characteristics (e.g., length and personality), the need to address the variance has been raised because it can lead to notably higher error rates for specific groups. To address this problem, we proposed a new method for group-specific threshold optimization. Unlike universal thresholding methods, which presume evenly distributed probability regardless of the stylistic features of the text, our method effectively captures the variations of the probability distribution of AI text classifiers across different subgroups, leading to classification with a lower discrepancy.
The novelty of this work resides within introducing a group-adaptive, post-hoc threshold optimization framework that operates without retraining of the detectors in the given settings by leveraging fairness and performance metrics. To best of our knowledge, no prior research has systematically examined threshold optimization in the context of AI text detectors although some research investigated thresholding methods for probabilistic classifiers. FairOPT represents a significant step toward AI safety by providing an alternative to the common practice of applying a single fixed threshold (e.g., 0.5, 0.75 or 0.97151).
While our work focuses primarily on text classifiers, the FairOPT framework can naturally extend to other modalities where detecting AI-generated content is critical for safety and copyright (e.g., image, video, and audio), including deepfake video and audio detectors. This method will be highly beneficial in post-processing, guiding the model toward decreased discrepancy, contingent upon the specific requirements and deployment context. We can also apply this method to all other probabilistic classifications when subgroup discrepancy matters (e.g., recidivism risk prediction or hiring decisions) by learning tailored decision thresholds that substantially mitigate bias with negligible loss in overall accuracy. Furthermore, while our primary focus has been improving fairness in the post-training stage, we hypothesize that the proposed method can also be integrated as an in-training mechanism. Finally, other approaches for threshold optimization, including Bayesian methods and convex optimization, can also be studied in future research.
VI Limitations
While FairOPT demonstrably reduces subgroup discrepancy on our English‐text benchmarks, it relies on only two features (character‐count bins and model‐inferred Big Five traits). Other text attributes (e.g., topic, syntax, language) may induce additional biases that our current feature set does not capture. Moreover, the stylistic labels are derived from a downstream BERT‐based personality predictor, which itself may introduce cultural bias. We focus on relaxing demographic parity—extensions to other fairness definitions (equal opportunity or equalized odds) may require different optimization objectives or could be fundamentally incompatible.
VII Future Works
Although our experiments focused on English-language AI-generated text detectors, FairOPT is a modality-agnostic, subgroup-specific thresholding framework that readily extends to new deployment contexts and populations: for previously unseen subgroups, thresholds can be recalibrated from available data or warm-started from similar existing groups (e.g., AI-written text in schools, AI-generated images on a platform, or deepfakes on YouTube). Beyond text, FairOPT applies to image, video, and audio deepfake detection—any setting with probabilistic outputs that benefits from post-processing to ensure equitable performance—and more broadly to probabilistic classification tasks with subgroup disparities (e.g., recidivism risk prediction, hiring decisions), where learning tailored decision thresholds can mitigate bias with minimal impact on overall accuracy. We recognize that both theoretical analysis and user-centered case studies will be important for validating these extensions across diverse, real-world deployments.
VIII Impact Statement
This paper introduces novel theoretical frameworks and methodologies to enhance the robustness of AI-generated text detection systems. By extracting and using features for subgroups such as text length and stylistic traits, our approach addresses disparities in classification outcomes. This technology can play a crucial role in AI detection for identifying misinformation [8], safeguarding the integrity of publication and academic organizations [48, 39], and countering potential cybersecurity threats [49]. Since it is crucial for these AI-content detection methods to be robust and fair across many potential users, our method takes a major step in this direction by formulating the problem setting and developing a new algorithm.
Acknowledgments
Thanks to Hengzhi Li, Megan Tjandrasuwita, David Dai, Jeongmin Kwon, Chanakya Ekbote, and Seonghyun Nam for constructive feedback and discussion.